adding and subtracting rational expressions

8/(x^2-4x) + 2/x
1
Expert's answer
2012-12-03T09:05:01-0500
8/(x^2-4x) + 2/x=8/((x-4)*x) + 2/x=8/((x-4)*x) +2(x-4)/[x(x-4)]=[8+2x-8]/[x(x-4)]=2x/(x*(x-4))=2/(x-4)